What quantity of 65 percent acid solution must be mixed with a 20 percent solution to produce 270 mL of a 50 percent solution?
Answer by robertb(5830)

0.65x + 0.20(270 - x) = 0.50*270
==> 0.65x + 54 - 0.20x = 135
==> 0.45x = 81
==> x = 180 ml, the amount of 65% acid , and
270 - x = 90 ml, the amount of 20% acid.